please help me solve this equation and please show how you did it - I keep getting
x=1 for the answer and I know it isn't correct.
-3(2x+9)=21

Found 2 solutions by tjnw79, mmm4444bot:
Answer by tjnw79(57)

-3(2x+9)=21
what you need to do first is times -3 by 2x and 9
(-3*2x)+(-3*9)=21
-6x+-27=21
then add 27 to both sides
-6x(+27)(-27)=21+(27)
-6x=48
divide -6 on both sides
x=-8

Answer by mmm4444bot(95)

It would help if you showed your work. Then I could see how you're getting x = 1.
The first step is to get rid of the parentheses. Use the distributive property. In other words, both the 2x and the 9 get multiplied by -3.
-6x - 27 = 21
Now, we want to isolate the term containing the variable. In other words, we want to get rid of the -27 on the left wide of the equals sign. We do this by adding 27 to both sides.
-6x = 48
The last step solves for x. Divide both sides by -6.
x = -8
Let's check our result by substituting our answer into the original equation to see if works.
-3[2(-8) + 9] = 21
-3[-16 + 9] = 21
-3[-7] = 21
21 = 21.
~ Mark